OpenStudy (anonymous):

Use continuity to evaluate the limit. As x approaches 9, what is the limit of [55+ (squareroot of x)]/ (squareroot of 55 +x)?

OpenStudy (anonymous):

\[55+ \sqrt{x}/ \sqrt{55+x}\]

OpenStudy (anonymous):

as x approaches 9, the limit is defined, you can just plug it in from what I see, if I don't misinterpret the equation

OpenStudy (anonymous):

I know that first I need to rationalize the denominator, which is what I'm having difficulty with. I end up with \[55+\sqrt{9}*\sqrt{55+9}/64\]

OpenStudy (anonymous):

I know that the correct answer is 52/8, but I am having trouble with the intermediate step.

OpenStudy (anonymous):

Is the equation \[ \large 55 + \frac{\sqrt{x}}{\sqrt{55+x}}\] or \[ \frac{55+\sqrt{x}}{\sqrt{55+x}} \]